OpenStudy (anonymous):

Which of the following is the smallest volume? (2 points) 500 mL 2,500 cm3 5.5 • 10–1 L 25 m3

OpenStudy (anonymous):

please help!!!

OpenStudy (anonymous):

You'll need your SI prefix conversion ability, plus the conversion between cm^3 and mL (1 cm^3 = 1 mL) to put them all in the same unit. Let's pick mL. 1) 500 mL = 500 mL. That was easy! 2) 2500 cm^3 = 2500 mL (using 1 cm^3 = 1 mL). 3) 5.5 x 10^-1 L = 0.55 L = 0.55 L * (1000 mL/L) = 550 mL 4) 25 m^3 = 25 m^3 (100 cm/m)^3 = 25 x 10^6 cm^3 = 25 x 10^6 mL. Now you can choose pretty easily.

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

What's with the question mark? If you're not sure, why not? What will make you sure? There's nothing wrong with being unsure -- it's a very scientific attitude, really -- but it's not possible to help yourself, or be helped, unless you can zero in on why you're unsure. Why are you unsure?

OpenStudy (anonymous):

its A im sure lol

OpenStudy (anonymous):

Hmm...well, OK. But please keep in mind education in science is not (or should not) be about just repeating the Accepted Wisdom, the truths and social myths we all "just know." To some extent it's about sharing with you the techniques (like SI prefixes) and discoveries of those who came before you. But the only rational reason for requiring education in science for those who aren't going to be actual scientists (which is nearly everybody) is if it has benefits outside of science. It does. What it does is give you a way of sharpening your natural skepticism. When anybody, in any area, tells you X and asks you to believe it, you will naturally be skeptical, to some degree or other. A good education in science teaches you how to sharpen that skepticism, to turn it into specific questions and challenges that help you better evaluate the truth of what you're being asked to accept. So be not afraid of "stupid" questions, or a lack of understanding, or unsureness. Those are things to be cherished, in science. The point is to learn to sharpen them to specific questions and experiments, if only mental experiments. "I don't understand your answer" is a feeling you may have. Where science helps you out is teaching you to focus that: "I don't understand your answer...because I don't see why you did what you did in line 2 of your calculation....because I don't see why you raised that conversion factor to the 3rd power...because the fact that you used "centi" on the left and "milli" on the right makes me question that equation...because I don't see how you can measure volume in the same units you use to measure length..." And so on. Once you sharpen doubt to a series of specific questions and challenges, you are about 90% on the way to resolving the doubt (into belief or unbelief), because you can get those questions answered.

OpenStudy (anonymous):

wow that was like an essay ^